STUDY ON THE SPECTROSCOPY OF TRANSITION METAL FLUORIDES

\underline{CONG-XIANG CHEN}, XIAO-PENG ZHANG, JIN JIN, QIN RAN, YANG CHEN, Laboratory of Bond Selective Chemistry, and Department of Chemical Phys ics, University of Science and Technology of China, Hefei, Anhui, 230026 P. R. China.

We present the spectra of NiF and CuF by using pulsed dc discharge to produce the species and laser induced fluorescence technique under free jet-cooled conditions. For NiF molecule, twenty one electro nic transition in the 400-580 nm region, involving five high-lying electronic states[18.1]^{2}Δ_{5/2}, [19.7]^{2}Π_{3/2},[20.0]^{2}Δ_{5/2}, [20.3]Π_{3/2} and [22.9] ^{2}Π_{i} and two lower states, the ground state X ^{2}Π_{3/2} and the low-lying state [0.25]^{2}Σ, respectively, are rotationally analyzed. The molecular constants of these five upper states are determined. The [20.3] Π_{3/2} state is identified and may be linked to the component of a unreported quartet state. A previously reported [20.0]^{2}Π_{1/2} component may be rea ssigned as a ^{2}Δ_{5/2}.

For CuF molecule, the spectra in the 450-510nm region are observed. They mainly result from the B-X and C-X transitions. The rotationally resolved bands B(^{1}Σ^{+}, v′)(v′ = 0 − 3) − X(^{1}Σ^{+}, v′′ = 0) and C(^{1}Π, v′)(v′=0-2)-X(^{1}Σ^{+}) are observed. The spectroscopic isotopic shifts between ^{65}CuF and ^{63}CuF are observed. The rotational constants for low and upper states are derived.